OpenStudy (anonymous):

convert each equation to standard form by completing the square on x and y. Then graph the ellipse and give the location of its foci. 4x^2+y^2+16x-6y-39=0 4x^2+25y^2-24x+100y+36=0

OpenStudy (anonymous):

and also (x+2)^2/16 + (y-3)^2 graph each elipse and give the location of its foci

OpenStudy (anonymous):

\[(x+2)^{2}\div16 + (y-3)^2=1\]

OpenStudy (anonymous):

i nned help on this one^

OpenStudy (anonymous):

4x^2+16x+y^2-6y-39=0 4(x^2+4x)+y^2-6y-39=0 4(x^2+4x+4-4)+(y^2-6y+9-9)-39=0 4(x+2)^2-16+(y-3)^2-9-39=0 4(x+2)^2+(y-3)^2=64 \[\frac{(x+2)^2}{16}+\frac{(y-3)^2}{64}=1\]

OpenStudy (anonymous):

center is (-2,3) vertical axis is the major axis c^2=64-16=48 c=4sqrt(3) focii(-2, 3 +/- 4sqrt(3))

OpenStudy (anonymous):

i need steps how do youdo it

OpenStudy (anonymous):

do what?

OpenStudy (anonymous):

how did you transfer the equation

OpenStudy (anonymous):

i showed all the steps, completed the square, put all the constants on the right, then divided by the constant

OpenStudy (anonymous):

-16-9-39=-64, add 64 to both sides

OpenStudy (anonymous):

nooo i need help on (x+2)^2/16 + (y-3)^2

OpenStudy (anonymous):

where'd that one come from?

OpenStudy (anonymous):

i need help on that one if you could

OpenStudy (anonymous):

does it equal one?

OpenStudy (anonymous):

my equation came from completing the square of the first equation in your question

OpenStudy (anonymous):

yess the equation = 1

OpenStudy (anonymous):